OpenStudy (anonymous):

Perform the operation. 3 ( a^2 b ) + 2 / ( ab^2 ) ~I keep getting 3/(a^b) + 2/(ab^2) and it isnt an answer~

OpenStudy (saifoo.khan):

\[3 ( a^2 b ) + \frac{2 }{ ( ab^2 ) }\]Like that?

OpenStudy (anonymous):

3 divided by (a^2b) plus 2 over (ab^2)

OpenStudy (anonymous):

\[\frac{3}{ a^2 b } + \frac{2 }{ ab^2 } \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

least common multiple of the denominators is \(a^2b^2\)

OpenStudy (anonymous):

\(\frac{3}{a^2b}=\frac{3b}{a^2b^2}\) and \(\frac{2}{ab^2}=\frac{2a}{a^2b^2}\) now you can add them

OpenStudy (anonymous):

i.e. put them over the same denominator and get \[\frac{2a+3b}{a^2b^2}\]

OpenStudy (anonymous):

thanks! i dont know why i got the other answer.